# Exoplanet Detection : A Detailed Analysis

### v1

@@ Abstract
Exoplanet Detection, Transit, Missions, Exoplanets, Gaia
@@ 1 Introduction
Since the discovery of exoplanets began, they have continually surprised us with the hidden secrets of the cosmos. Here on earth, wealth is often associated with the possession of precious rocks and metals. But what about worlds made of these rocks and metals? For instance, 55 Cancri e is an exoplanet made of diamonds and orbits a sun-like star. It was discovered by McArthur McArthur et al. (2004) using the radial velocity method by observing the spectrum of one of the parent stars in a binary system located 40 light-years from the planet Earth. Exoplanets are poised to become a significant part of human civilization in various aspects, such as energy sources, colonization of other planets, transportation, space exploration, employment, and many others. Since the first exoplanet was discovered, astronomers have been working tirelessly to discover other extrasolar planets. Numerous missions have been conducted by several space agencies to study exoplanets, and the results have been truly fulfilling. Till date, approximately 5500+ exoplanets have been discovered Akeson et al. (2013), with many of them undergoing detailed studies.
@@
51 Pegasi b, the first exoplanet, orbiting around a star similar to sun, was found using the radial velocity approach by Mayor & Queloz. Several space agencies are running missions specifically for exoplanet detection using different detection methods Shafi (2021). These missions include hubble, spitzer, corot, kepler/K2, gaia, TESS, CHEOPs, webb, and many more, which continuously monitor the universe, and astronomers keep detecting exoplanets with their help. In the future, there will be missions for specific exoplanets to understand their properties in detail and their environmental conditions. However, as of now, it’s hard to travel to even nearby exoplanets within a lifetime. Further more, introduced space missions briefly, including details about data formats and other details. Astronomers use various machine learning Kumari et al. (2023) and data cleaning algorithms to work on the data that comes from these missions. Such algorithms include the BLS algorithm, light curve analyzer, CNN, and other algorithms. all exoplanets are formed by normal matter but some scientists assumes that some exoplanets could me formed through dark matter Bai et al. (2023). Introduction about exoplanets with there basic details are going to be dicsussed in the next section.
@@ 2 The exoplanets
Extrasolar planets, also referred to as exoplanets, are planets that circle stars outside of our solar system. From the first exoplanet to be discovered in the early 1990s to the current booming field with thousands of confirmed planets, the growth of exoplanet investigation has been incredible wikipedia (2022). Section 2 discusses the crucial roles that various detection methods have had in this process. Technological developments like adaptive optics, space-based observatories like kepler and TESS, and high-precision spectrographs have made these approaches far more useful Woodrum et al. (2023). The discovery of smaller and farther-off planets has been made feasible by these developments. Exoplanet research is essential because it can increase our knowledge of planetary systems and the conditions necessary for habitability and the possibility that life exists elsewhere in the cosmos.
@@ 2.1 Stars That Host Planets
On average, each star Scalo et al. (2007) is believed to host at least one planet. It’s estimated that around 20% of sun like stars have an "earth-sized" planet in their habitable zones. The majority of known exoplanets orbit stars similar to our Sun, specifically those in the main-sequence spectral categories F, G, or K Lee et al. (2012) Perryman (2018) Meunier et al. (2023). Planets detectable by the radial-velocity method are less common around lower mass stars, such as the red dwarfs in the M spectral group Ananyeva et al. (2023). Despite this, the transit method, which excels at detecting smaller planets, has enabled the kepler telescope to identify dozens of planets orbiting near red dwarfs.
@@ 2.2 Detection methods
Various detection methods Rouan & Lagrange (2023) exist for exoplanets, and astronomers continuously work toward developing new ways of detection. Some effective methods include radial velocity, which led to the discovery of the 1st exoplanet in the late ’90s; the transit method, which analyzes light curves and periodograms Bozza et al. (2016); direct imaging, capturing exoplanets through images; gravitational microlensing; and astrometry Perryman (2018). Table 1 shows the total number of exoplanets discovered by various methods of detection till beginning of 2024 NASA (2024). data taken from NASA exoplanet archive Akeson et al. (2013).
@@
Discovered Exoplanets
Method of discovery Number of planets
Astrometry 3
@@ 2.3 Types of exoplanets
As of the conclusion of 2023, the existence of over 5500 planets had been confirmed. In an attempt to better understand these celestial bodies, scientists have undertaken numerous initiatives to categorize them based on various attributes, including their size, mass, and composition Perryman (2018).Table 2 shows the classification of exoplanets based on mass, size and chemical compositions of known gases such as $ZnS$, $H_{2}O$, $CO_{2}$ $CH_{4}$ Forestano et al. (2023).
Figure 3: planet masses are categorized based on solar system objects, as suggested by Stevens & Gaudi (2013). The solar system’s planet masses are displayed on the right, with the size of the circles corresponding to the planets’ radii Perryman (2018) Mathur et al. (2021).
@@
Factor Planet type Parameter
Size Earth size ($1.25-2R_{\oplus}$)
@@ 2.4 Atmospheric conditions
Finding transiting exoplanets is a significant breakthrough that provides previously unattainable knowledge about the composition and physical properties of these far-off worlds Yakovlev et al. (2022). Transit and eclipse spectroscopy allows the investigation of atmospheric dynamics Walker et al. (2018) and heat transport processes, especially for hot Jupiters, while the resulting densities, derived from detailed measurements of masses and radii, offer important hints regarding internal structures and compositions. These discoveries have been made possible in large part by drawing on knowledge amassed over the previous 50 years from research on the planets and satellites of the solar system Jakka (2023).
Future developments in photometric and spectroscopic observations, such as those brought about by the 2020 launch of the James Webb Space Telescope (JWST), are very promising. Advanced diagnostics may reveal the precise atmospheric compositions of exoplanets that are located in the "habitable zone Jagtap et al. (2021), " which is where liquid water may be present. These compositions may include indications of either primitive or advanced life forms. These developments advance not only the hunt for extraterrestrial life but also help answer the important question of how common or uncommon life is in the universe Seager & Deming (2010). The dynamic interplay between theoretical models and observational methods is set to propel future research and transform our knowledge of exoplanetary systems.
@@ 3 Exoplanet detection methods
As discussed previously there are various detection methods which played an important role to present this unique field of study. all methods works on various different properties such as light, noise Smith et al. (2008), EMwaves etc. some of them are explained here.
@@ 3.1 Radial Velocity Method
Radial Velocity: When a planet orbits a star, both celestial bodies revolve around a shared center of mass, termed the barycenter. The star’s motion around this point results in a periodic fluctuation in its radial velocity Hatzes (2016), which signifies its speed moving towards or away from an observer located on Earth. It is discovered by spectroscopic studies that this radial velocity shift exists. The spectral lines of a star change towards the blue end of the spectrum as it travels towards the observer, and to the red end when it goes away (blueshift and redshift, respectively) Bozza et al. (2016) Chen (2023).
Angular (Astrometric) Position in the Sky: The star moves in a modest circular or elliptical path around the shared center of mass due to the gravitational attraction of an orbiting planet Pham et al. (2024) Bozza et al. (2016). The star’s angular location in the sky changes periodically, indicating its motion. Astrometric measures entail monitoring the star’s apparent location with accuracy over time. Measurements of radial velocity have been used more frequently than this approach, which is more difficult Perryman (2018) Chen (2023).
Time of Arrival of Periodic Reference Signal: Variations in the arrival times of periodic signals, such pulsar pulses or eclipse timings in binary star systems, can be caused by the gravitational interaction between a star and its circling planet Bozza et al. (2016). Timing variations in these signals of reference may indicate the existence of an unseen planet and offer additional evidence of its impact Fischer et al. (2015).
Radial velocity Dumusque et al. (2017) measurements revealed the first credible discovery of an exoplanet in 1995 Perryman (2018). Since then, a number of exoplanets have been found using this method because to technological and observational breakthroughs. Via radial velocity metrics, 1075 exoplanets had been found as of 02/05/2024; some of these planets have been observed to transit their host stars, allowing the star’s light to periodically fade as the planet passes in front of it.
@@ 3.2 Transit method
Astronomy achieved a major milestone in 1999 with the ground-breaking observation of the first transit of an extrasolar planet Perryman (2018) Chen (2023). The ’hot Jupiter’ system HD 209458 was observed photometrically Perryman (2018), which allowed for this observation. A gas giant planet known as a "hot Jupiter" has a close orbit around its host star, which increases the planet’s temperature somewhat. other star systems found by radial velocity surveys came into focus after this finding Fischer et al. (2015). After a planet passes in front of its host star, causing a periodic dimming of the star’s brightness, researchers kept an eye on these systems in the hopes of finding other transits Yang (2023). Identifying exoplanets became increasingly dependent on this methodology, particularly for those that were not initially found via the radial velocity method Bozza et al. (2016). Dedicated surveys were rapidly launched to perform "blind hunts," initially from space then from ground-based observatories Perryman (2018). The goal of these surveys was to find new planets only by using the periodic transit fingerprints observed in star light curves. With this method, scientists were able to find a wide variety of exoplanets without any prior awareness of their presence Chen (2023).
@@
In the field of exoplanet research, transiting planets are especially significant Afanasev (2018). The light curves acquired during transits yield important details on planet sizes since the radii of the planets may be estimated due to the star’s periodic dimming of light Perryman (2018) Fischer et al. (2015). Through techniques like radial velocity measurements, astronomers can establish the transiting planet’s mass as well as its density, providing preliminary information about the planet’s composition Bozza et al. (2016) Chen (2023).
@@
Around the end of the year of 2023, exoplanet research had undergone a major transformation. There were now close to 4153 transiting planets known, with the Kepler space observatory detecting nearly 2778 of them as reported by NASA. Kepler began operations in 2009 and has been a key player in the area of transiting exoplanet research, providing an abundance of data that has made significant effects Panahi et al. (2022) Chen (2023). A wide range of exoplanetary systems have been found as a result of the joint efforts of ground-based surveys Perryman (2018), such HAT and WASP, which revealed bright and easily detectable transiting systems, and Kepler’s numerous discoveries, particularly those of close-orbiting multi-planet systems Afanasev (2018). These systems contribute to our understanding of the diversity of planetary systems outside of our solar system by displaying a wide range of physical attributes and circumstances.
@@ 3.3 Direct imaging method
"Imaging" in the context of modern exoplanet research is the process of recognizing a planet as an individual source of light Perryman (2018) Chen (2023). This may happen as a result of the planet’s own thermal emission or light reflected from the host star. It is important that this imaging idea is not the same as reaching the planetary surface’s spatial resolution Bozza et al. (2016). Additionally, it is distinct from techniques that detect fluctuations in the total brightness Manfredi et al. (2020) of the planet and star system as a result of things like secondary eclipses, reflected light, or changes in the stellar spectrum Birkby (2018) throughout a transit. In 2005, the first imaging detections were reported, namely the identification of huge, youthful, self-luminous planets in broad orbits Bozza et al. (2016) Perryman (2018). The NASA exoplanet archive listed 44 of these discoveries as at the end of 2017 Perryman (2018). Prominent instances comprise planets discovered in well researched debris disk systems such as $\beta$ Pic and fomalhaut, in along with multi-planet systems like HR 8799 (which includes four planets) and LkCa 15 (containing two planets) Perryman (2018). There are certain deep learning models are used to inhance the capabilities of detection such as deep PACO Flasseur et al. (2024)
Researchers have been able to determine the orbital velocity of planets in systems such as $\beta$ Pic, HR 8799, and fomalhaut thanks to multi-epoch research. Over the past ten years, developments in imaging instrument technology Daglayan et al. (2023) Hebrard et al. (2018) have led to an emergence of cutting-edge devices specifically intended for exoplanet imaging Beuzit et al. (2006). VLT-SPHERE, Gemini-GPI Galicher et al. (2014), and Subaru-HiCIAO are a few examples Fischer et al. (2015). It is important to remember that these devices can only directly picture massive planets that are young, warm, and self-luminous at this time Perryman (2018). It is expected that with the next 30–40-meter telescopes, direct detection of point source pictures resulting from reflected starlight would become possible Bozza et al. (2016) Perryman (2018). On the other hand, picturing planets similar to earth is still a distant objective Chen (2023). It is crucial to take into account imaging Koechlin et al. (2005) at radio and X-ray wavelengths while studying star-planet systems. These other wavelength ranges may provide light on many facets of the interactions that take place between stars and their planets Flasseur et al. (2023). There is talk about the potential to expand imaging capabilities to X-ray and radio wavelengths, which implies a more thorough investigation of observational methods in exoplanet studies Rosa & do Nascimento Dias (2023) Daglayan et al. (2023).

### v2

@@ Abstract
Exoplanet Detection—Transit—Missions—Exoplanets—Gaia.
[…]
mahimakaushik333@gmail.com
[…]
@@ 1 Introduction
Since the discovery of exoplanets began, they have continually surprised us with the hidden secrets of the cosmos. Here on earth, wealth is often associated with the possession of precious rocks and metals. But what about worlds made of these rocks and metals? For instance, 55 Cancri e is an exoplanet made of diamonds and orbits a sun-like star. It was discovered by McArthur McArthur $et~{}al$. (2004) using the radial velocity method by observing the spectrum of one of the parent stars in a binary system located 40 light-years from the planet Earth. Exoplanets are poised to become a significant part of human civilization in various aspects, such as energy sources, colonization of other planets, transportation, space exploration, employment, and many others. Since the first exoplanet was discovered, astronomers have been working tirelessly to discover other extrasolar planets. Numerous missions have been conducted by several space agencies to study exoplanets, and the results have been truly fulfilling. Till date, approximately 5500+ exoplanets have been discovered Akeson $et~{}al$. (2013), with many of them undergoing detailed studies.
@@
51 Pegasi b, the first exoplanet, orbiting around a star similar to sun, was found using the radial velocity approach by Mayor & Queloz. Several space agencies are running missions specifically for exoplanet detection using different detection methods Shafi (2021). These missions include hubble, spitzer, corot, kepler/K2, gaia, TESS, CHEOPs, webb, and many more, which continuously monitor the universe, and astronomers keep detecting exoplanets with their help. In the future, there will be missions for specific exoplanets to understand their properties in detail and their environmental conditions. However, as of now, it’s hard to travel to even nearby exoplanets within a lifetime. Further more, introduced space missions briefly, including details about data formats and other details. Astronomers use various machine learning Kumari $et~{}al$. (2023) and data cleaning algorithms to work on the data that comes from these missions. Such algorithms include the BLS algorithm, light curve analyzer, CNN, and other algorithms. all exoplanets are formed by normal matter but some scientists assumes that some exoplanets could me formed through dark matter Bai $et~{}al$. (2023). Introduction about exoplanets with there basic details are going to be dicsussed in the next section.
@@ 2 The exoplanets
Extrasolar planets, also referred to as exoplanets, are planets that circle stars outside of our solar system. From the first exoplanet to be discovered in the early 1990s to the current booming field with thousands of confirmed planets, the growth of exoplanet investigation has been incredible wikipedia (2022). Section 2 discusses the crucial roles that various detection methods have had in this process. Technological developments like adaptive optics, space-based observatories like kepler and TESS, and high-precision spectrographs have made these approaches far more useful Woodrum $et~{}al$. (2023). The discovery of smaller and farther-off planets has been made feasible by these developments. Exoplanet research is essential because it can increase our knowledge of planetary systems and the conditions necessary for habitability and the possibility that life exists elsewhere in the cosmos.
@@ 2.1 Stars That Host Planets
On average, each star Scalo $et~{}al$. (2007) is believed to host at least one planet. It’s estimated that around 20% of sun like stars have an ”earth-sized” planet in their habitable zones. The majority of known exoplanets orbit stars similar to our Sun, specifically those in the main-sequence spectral categories F, G, or K Lee $et~{}al$. (2012) Perryman (2018) Meunier $et~{}al$. (2023). Planets detectable by the radial-velocity method are less common around lower mass stars, such as the red dwarfs in the M spectral group Ananyeva $et~{}al$. (2023). Despite this, the transit method, which excels at detecting smaller planets, has enabled the kepler telescope to identify dozens of planets orbiting near red dwarfs.
@@ 2.2 Detection methods
Various detection methods Rouan & Lagrange (2023) exist for exoplanets, and astronomers continuously work toward developing new ways of detection. Some effective methods include radial velocity, which led to the discovery of the 1st exoplanet in the late ’90s; the transit method, which analyzes light curves and periodograms Bozza $et~{}al$. (2016); direct imaging, capturing exoplanets through images; gravitational microlensing; and astrometry Perryman (2018). Table 1 shows the total number of exoplanets discovered by various methods of detection till beginning of 2024 NASA (2024). data taken from NASA exoplanet archive Akeson $et~{}al$. (2013).
[…]
@@
[…]
@@ 2.3 Types of exoplanets
As of the conclusion of 2023, the existence of over 5500 planets had been confirmed. In an attempt to better understand these celestial bodies, scientists have undertaken numerous initiatives to categorize them based on various attributes, including their size, mass, and composition Perryman (2018).Table 2 shows the classification of exoplanets based on mass, size and chemical compositions of known gases such as $ZnS$, $H_{2}O$, $CO_{2}$ $CH_{4}$ Forestano $et~{}al$. (2023).
Figure 3: planet masses are categorized based on solar system objects, as suggested by Stevens & Gaudi (2013). The solar system’s planet masses are displayed on the right, with the size of the circles corresponding to the planets’ radii Perryman (2018) Mathur $et~{}al$. (2021).
[…]
@@
[…]
@@ 2.4 Atmospheric conditions
Finding transiting exoplanets is a significant breakthrough that provides previously unattainable knowledge about the composition and physical properties of these far-off worlds Yakovlev $et~{}al$. (2022). Transit and eclipse spectroscopy allows the investigation of atmospheric dynamics Walker $et~{}al$. (2018) and heat transport processes, especially for hot Jupiters, while the resulting densities, derived from detailed measurements of masses and radii, offer important hints regarding internal structures and compositions. These discoveries have been made possible in large part by drawing on knowledge amassed over the previous 50 years from research on the planets and satellites of the solar system Jakka (2023).
Future developments in photometric and spectroscopic observations, such as those brought about by the 2020 launch of the James Webb Space Telescope (JWST), are very promising. Advanced diagnostics may reveal the precise atmospheric compositions of exoplanets that are located in the ”habitable zone Jagtap $et~{}al$. (2021), ” which is where liquid water may be present. These compositions may include indications of either primitive or advanced life forms. These developments advance not only the hunt for extraterrestrial life but also help answer the important question of how common or uncommon life is in the universe Seager & Deming (2010). The dynamic interplay between theoretical models and observational methods is set to propel future research and transform our knowledge of exoplanetary systems.
@@ 3 Exoplanet detection methods
As discussed previously there are various detection methods which played an important role to present this unique field of study. all methods works on various different properties such as light, noise Smith $et~{}al$. (2008), EMwaves etc. some of them are explained here.
@@ 3.1 Radial Velocity Method
Radial Velocity: When a planet orbits a star, both celestial bodies revolve around a shared center of mass, termed the barycenter. The star’s motion around this point results in a periodic fluctuation in its radial velocity Hatzes (2016), which signifies its speed moving towards or away from an observer located on Earth. It is discovered by spectroscopic studies that this radial velocity shift exists. The spectral lines of a star change towards the blue end of the spectrum as it travels towards the observer, and to the red end when it goes away (blueshift and redshift, respectively) Bozza $et~{}al$. (2016) Chen (2023).
Angular (Astrometric) Position in the Sky: The star moves in a modest circular or elliptical path around the shared center of mass due to the gravitational attraction of an orbiting planet Pham $et~{}al$. (2024) Bozza $et~{}al$. (2016). The star’s angular location in the sky changes periodically, indicating its motion. Astrometric measures entail monitoring the star’s apparent location with accuracy over time. Measurements of radial velocity have been used more frequently than this approach, which is more difficult Perryman (2018) Chen (2023).
Time of Arrival of Periodic Reference Signal: Variations in the arrival times of periodic signals, such pulsar pulses or eclipse timings in binary star systems, can be caused by the gravitational interaction between a star and its circling planet Bozza $et~{}al$. (2016). Timing variations in these signals of reference may indicate the existence of an unseen planet and offer additional evidence of its impact Fischer $et~{}al$. (2015).
Radial velocity Dumusque $et~{}al$. (2017) measurements revealed the first credible discovery of an exoplanet in 1995 Perryman (2018). Since then, a number of exoplanets have been found using this method because to technological and observational breakthroughs. Via radial velocity metrics, 1075 exoplanets had been found as of 02/05/2024; some of these planets have been observed to transit their host stars, allowing the star’s light to periodically fade as the planet passes in front of it.
@@ 3.2 Transit method
Astronomy achieved a major milestone in 1999 with the ground-breaking observation of the first transit of an extrasolar planet Perryman (2018) Chen (2023). The ’hot Jupiter’ system HD 209458 was observed photometrically Perryman (2018), which allowed for this observation. A gas giant planet known as a ”hot Jupiter” has a close orbit around its host star, which increases the planet’s temperature somewhat. other star systems found by radial velocity surveys came into focus after this finding Fischer $et~{}al$. (2015). After a planet passes in front of its host star, causing a periodic dimming of the star’s brightness, researchers kept an eye on these systems in the hopes of finding other transits Yang (2023). Identifying exoplanets became increasingly dependent on this methodology, particularly for those that were not initially found via the radial velocity method Bozza $et~{}al$. (2016). Dedicated surveys were rapidly launched to perform ”blind hunts,” initially from space then from ground-based observatories Perryman (2018). The goal of these surveys was to find new planets only by using the periodic transit fingerprints observed in star light curves. With this method, scientists were able to find a wide variety of exoplanets without any prior awareness of their presence Chen (2023).
@@
In the field of exoplanet research, transiting planets are especially significant Afanasev (2018). The light curves acquired during transits yield important details on planet sizes since the radii of the planets may be estimated due to the star’s periodic dimming of light Perryman (2018) Fischer $et~{}al$. (2015). Through techniques like radial velocity measurements, astronomers can establish the transiting planet’s mass as well as its density, providing preliminary information about the planet’s composition Bozza $et~{}al$. (2016) Chen (2023).
@@
Around the end of the year of 2023, exoplanet research had undergone a major transformation. There were now close to 4153 transiting planets known, with the Kepler space observatory detecting nearly 2778 of them as reported by NASA. Kepler began operations in 2009 and has been a key player in the area of transiting exoplanet research, providing an abundance of data that has made significant effects Panahi $et~{}al$. (2022) Chen (2023). A wide range of exoplanetary systems have been found as a result of the joint efforts of ground-based surveys Perryman (2018), such HAT and WASP, which revealed bright and easily detectable transiting systems, and Kepler’s numerous discoveries, particularly those of close-orbiting multi-planet systems Afanasev (2018). These systems contribute to our understanding of the diversity of planetary systems outside of our solar system by displaying a wide range of physical attributes and circumstances.
@@ 3.3 Direct imaging method
”Imaging” in the context of modern exoplanet research is the process of recognizing a planet as an individual source of light Perryman (2018) Chen (2023). This may happen as a result of the planet’s own thermal emission or light reflected from the host star. It is important that this imaging idea is not the same as reaching the planetary surface’s spatial resolution Bozza $et~{}al$. (2016). Additionally, it is distinct from techniques that detect fluctuations in the total brightness Manfredi $et~{}al$. (2020) of the planet and star system as a result of things like secondary eclipses, reflected light, or changes in the stellar spectrum Birkby (2018) throughout a transit. In 2005, the first imaging detections were reported, namely the identification of huge, youthful, self-luminous planets in broad orbits Bozza $et~{}al$. (2016) Perryman (2018). The NASA exoplanet archive listed 44 of these discoveries as at the end of 2017 Perryman (2018). Prominent instances comprise planets discovered in well researched debris disk systems such as $\beta$ Pic and fomalhaut, in along with multi-planet systems like HR 8799 (which includes four planets) and LkCa 15 (containing two planets) Perryman (2018). There are certain deep learning models are used to inhance the capabilities of detection such as deep PACO Flasseur $et~{}al$. (2024)
Researchers have been able to determine the orbital velocity of planets in systems such as $\beta$ Pic, HR 8799, and fomalhaut thanks to multi-epoch research. Over the past ten years, developments in imaging instrument technology Daglayan $et~{}al$. (2023) Hebrard $et~{}al$. (2018) have led to an emergence of cutting-edge devices specifically intended for exoplanet imaging Beuzit $et~{}al$. (2006). VLT-SPHERE, Gemini-GPI Galicher $et~{}al$. (2014), and Subaru-HiCIAO are a few examples Fischer $et~{}al$. (2015). It is important to remember that these devices can only directly picture massive planets that are young, warm, and self-luminous at this time Perryman (2018). It is expected that with the next 30–40-meter telescopes, direct detection of point source pictures resulting from reflected starlight would become possible Bozza $et~{}al$. (2016) Perryman (2018). On the other hand, picturing planets similar to earth is still a distant objective Chen (2023). It is crucial to take into account imaging Koechlin $et~{}al$. (2005) at radio and X-ray wavelengths while studying star-planet systems. These other wavelength ranges may provide light on many facets of the interactions that take place between stars and their planets Flasseur $et~{}al$. (2023). There is talk about the potential to expand imaging capabilities to X-ray and radio wavelengths, which implies a more thorough investigation of observational methods in exoplanet studies Rosa & do Nascimento Dias (2023) Daglayan $et~{}al$. (2023).